# Physics of droplet regulation in biological cells

### v1

@@ 2.1.4  Interfaces and phase morphologies
where $\bar{\mu}_{\mathrm{eq}}$ is the exchange chemical potential and $\Pi_{\mathrm{eq}}=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{out}}-f(\phi^{(0)}_{\mathrm{out}})=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{in}}-f(\phi^{(0)}_{\mathrm{in}})$ denotes the associated osmotic pressure; see appendix C. The second line assumes a monotonic profile $\phi_{\mathrm{eq}}(x)$, which might not be valid in complex mixtures [128]. However, $\phi_{\mathrm{eq}}(x)$ is always monotonic in binary mixtures, and we find in the symmetric case ($\Omega=1$)
@@
The thin-interface approximation simplifies the analysis of equilibrium states in binary liquids: material conservation fixes the total volume of the dilute and dense phases, so their shape is the only degree of freedom. The total free energy thus comprises a constant term covering the bulk contributions and an interfacial term, $\gamma A$, where $A$ is the total interfacial area. Equilibrium states thus minimise $A$, which implies zero mean curvature of the interfaces [131]. Consequently, surface energies can be interpreted as a mechanical surface tension [132], explaining why the minority phase favours spherical shapes. Conversely, if both phases occupy similar volumes, bicontinuous structures, such as cylinders and flat lamella, can emerge; see figure 2.2. These structures predominate in a region of phase space that overlaps significantly with the spinodal region; see figure 2.1E,F. However, these two phenomena are distinct, so droplets can for instance emerge from spinodal decomposition. Bicontinuous structures are rarely observed in biological cells, so we exclusively focus on droplets in this review.
@@ 2.2 Effective description of droplet dynamics
To focus on droplets, we next consider situations in which the dense phase forms compact regions separated from the dilute phase by a thin interface. For simplicity, we focus on three-dimensional droplets (key results for two dimensions appear in appendix D) without reactions (which are added in section 5).
@@ 2.2.1  Local equilibrium at the interface
where $V_{\mathrm{sys}}$ is the system volume. For simplicity, we consider a spherical droplet of radius $R$ (implying volume $V=\frac{4\pi}{3}R^{3}$ and surface area $A=4\pi R^{2}$) with a surface tension $\gamma$ that is independent of the interface curvature, but this approximation can in principle be lifted [133]. Material conservation implies
@@
where $\bar{\phi}$ is the average fraction in the system. Minimising $F$ using this constraint and fixed $V_{\mathrm{sys}}$ results in the coexistence conditions [23]
$\bar{\mu}(\phi_{\mathrm{in}})$ $=\bar{\mu}(\phi_{\mathrm{out}})$ (2.30a)
@@
where $d$ is the dimension of space. These equations reduce to equations (2.5) if surface tension is negligible, $\gamma=0$. However, the mechanical equilibrium given by equation (2.30b) reveals that $\gamma$ generates an extra pressure $P_{\gamma}=(d-1)\gamma/R$, which is known as the Laplace pressure. Equations (2.30) are solved by the coexisting equilibrium fractions $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$, which depend on the droplet radius $R$ and only coincide with $\phi^{(0)}_{\mathrm{in}}$ and $\phi^{(0)}_{\mathrm{out}}$ for flat interfaces ($R\rightarrow\infty$). A pressure increase $\Delta P$ inside droplets generally implies larger chemical potentials, and thus larger $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$. In the simple case of an incompressible droplet originating from strong phase separation, the fraction inside will hardly change, $\phi^{\mathrm{eq}}_{\mathrm{in}}\approx\phi^{(0)}_{\mathrm{in}}$, whereas the volume fraction outside approximately reads [134]
@@ 2.2.2  Classical nucleation theory
$J_{\mathrm{nucl}}$ $=J_{0}\exp\left(-\frac{\Delta\!F_{\mathrm{nucl}}}{k_{\mathrm{B}}T}\right)$ (2.35)
@@ 2.2.3  Growth of a single droplet
Once a droplet has nucleated, it can grow by taking up material from its surroundings or by creating material inside via chemical reactions. For simplicity, we postpone the discussion of reactions to section 5 and first focus on the case without reactions ($s=0$). Growth of droplets of any shape can be described by the normal velocity $v_{\perp}$ of the droplet interface, which is given by [139, 23]
@@
where $\boldsymbol{n}$ is the normal vector of the interface. Here, $\boldsymbol{j}_{\mathrm{in}}$ and $\boldsymbol{j}_{\mathrm{out}}$ denote the fluxes directly inside and outside of the interface, so the difference is related to the net accumulation of material. The fluxes follow from the Cahn–Hilliard equation (2.23), but the thin-interface limit also provides a useful simplification since the fractions $\phi$ typically only vary little inside and outside of droplets. Consequently, we linearize equation (2.23) around the compositions $\phi^{(0)}_{\mathrm{in}}$ and $\phi^{(0)}_{\mathrm{out}}$ to obtain a simple diffusion equation
@@
where we dropped the fourth-order derivative of $\phi$ and introduced the diffusivity
@@
As a simple example, we next consider a spherical droplet embedded in a large system. Without chemical reactions, the droplet is essentially homogeneous, implying $\boldsymbol{j}_{\mathrm{in}}=0$. Conversely, the fraction $\phi_{\infty}$ far away from the droplet might deviate from its equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$, resulting in diffusive fluxes around the droplet. Solving equation (2.37) in stationary state in three dimensions, we find $\boldsymbol{j}_{\mathrm{out}}=(\phi^{\mathrm{eq}}_{\mathrm{out}}-\phi_{\infty})D_{\mathrm{out}}R^{-1}$. The droplet growth rate following from equation (2.36) reads [140]
@@
where we neglected small surface tension effects in the denominator. Consequently, droplets grow spontaneously when the surrounding is supersaturated, $\phi_{\infty}>\phi^{\mathrm{eq}}_{\mathrm{out}}$. Moreover, droplets exhibit a drift when the surrounding concentration field is anisotropic [23],
@@
where the gradients are evaluated at the droplet position $\boldsymbol{x}$ and $d$ is the dimension of the system. Here, the anisotropy can either originate from concentration variations in the surrounding field $\phi_{\infty}$ or the equilibrium concentration itself might vary, e.g., due to gradients in additional, regulating components [141, 110] or mechanical properties [142, 143].
The drift caused by anisotropic fields is often dominated by other processes that affect droplets’ positions. For instance, droplets in cells can be actively moved by molecular motors [144] and hydrodynamic flows [110]. Moreover, all droplets exhibit Brownian motion characterised by a diffusion constant [145]
@@
where $\eta$ denotes the viscosity of the surrounding liquid, which might depend on the droplet’s size in complex liquids like the cytosol [146]. Equation (2.40) illustrates that droplets drift up supersaturation gradients, but Brownian motion caused by thermal fluctuations [145], hydrodynamic effects (discussed in section 3.3) and active transport can dominate their trajectory.
@@
which is an explicit equation for $V$ in the limit $\phi^{\mathrm{eq}}_{\mathrm{out}}\approx\phi^{(0)}_{\mathrm{out}}$. This equation illustrates that droplet size is controlled by material conservation described by equation (2.29), and that no droplets exist for concentrations left of the binodal line ($\bar{\phi}<\phi^{(0)}_{\mathrm{out}}$).
@@ 2.2.4  Dynamics of many droplets
The dynamics that we have derived for individual droplets allow us to discuss the interaction of many droplets as long as they are well-separated. We thus consider a collection of ${N_{\mathrm{d}}}$ droplets described by their positions $\boldsymbol{x}_{n}$ and radii $R_{n}$ for $n=1,\ldots,{N_{\mathrm{d}}}$. These droplets are embedded in a common dilute phase, whose composition is described by a field $\phi(\boldsymbol{x})$. In the typical situation where droplets occupy only a small fraction of the entire system, it is convenient to assume that $\phi(\boldsymbol{x})$ is defined in the entire system, and droplets merely are localised perturbations [140]. In this case, we describe the dynamics of the dilute phase by
@@
where the first term accounts for diffusive transport, while the second term accounts for the material exchange with the droplets of volume $V_{n}=\frac{4\pi}{3}R_{n}^{3}$ localised at $\boldsymbol{x}_{n}$. The growth and drift dynamics of the droplets follow from equations (2.39) and (2.40),
$\frac{\text{d}V_{n}}{\text{d}t}$ $=\frac{4\pi R_{n}D_{\mathrm{out}}}{\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}}\Bigl{(}\phi(\boldsymbol{x}_{n})-\phi^{\mathrm{eq}}_{\mathrm{out}}(R_{n})\Bigr{)}$ (2.44a)
@@
Equations (2.43)–(2.44) describe how droplets exchange material via diffusion through the dilute phase. In particular, droplets grow by taking up material from the dilute phase when the fraction $\phi$ in their surrounding is larger than the saturation concentration $\phi^{\mathrm{eq}}_{\mathrm{out}}$. Equation (2.32) shows that $\phi^{\mathrm{eq}}_{\mathrm{out}}$ decreases with the droplet radius $R$, implying larger droplets grow at the expense of smaller droplets. The resulting coarsening process, where the droplet count decreases while the mean size increases, is known as Ostwald ripening [147, 148]. In the limit where the diffusion of material between droplets is faster than their growth ($D_{\mathrm{out}}\gg L\,\text{d}R/\text{d}t$, where $L$ is the typical droplet separation), heterogeneities in the dilute phase equilibrate quickly, so all droplets exhibit a similar fraction $\phi(\boldsymbol{x})\approx\phi_{\infty}$. In this case, equation (2.32) implies that droplets grow when their radius is larger than the critical radius
@@
so $\langle V\rangle\propto t$ and $\langle{N_{\mathrm{d}}}\rangle\propto t^{-1}$ in $d=3$ dimensions.
Droplets also diffuse and coalesce, leading to larger and fewer droplets over time. Interestingly, this coarsening process exhibits the same scaling law as Ostwald ripening [151]. To see this, consider that coalescence roughly doubles the droplet volume, $\Delta\!V\propto R^{3}$, and the time between collisions scales as $\Delta t\propto L^{2}/D_{\mathrm{drop}}$ where $L\propto R$ is the average droplet separation and equation (2.41) implies $D_{\mathrm{drop}}\propto R^{-1}$. Taken together, $\Delta\!V/\Delta t\propto 1$, so the average volume grows linearly in time, similarly to Ostwald ripening. Although both coarsening processes obey the same scaling laws, the pre-factors can be very different and depend on particular details of the system [23].
Droplet coarsening is generally driven by the thermodynamic tendency to reduce interfacial area. However, the coarsening dynamics depend on details of kinetic processes [152], so observed exponents vary with composition and strength of phase separation [106]. For instance, coarsening can be slowed down by trapped species [153], charges [154], sub-diffusion in the dilute phase [155], and elastic inhomogeneities [156]. In contrast, coarsening can be accelerated when it is conversion-limited [150, 155], when mechanical relaxation is relevant [157], and when hydrodynamic effects are significant [158, 159], e.g., via coalescence-induced coalescence [160, 161]. In biological cells, coarsening is further complicated by active production and degradation of material, which we consider in section 5. It is currently unclear what governs the size distribution of droplets in cells, although there is evidence that a competition between coalescence and nucleation is crucial [162] and that the material properties of the surrounding are important [163]. Taken together, this shows that cells can use many physical processes to control droplet size.
@@ 2.3.3  Temporal variations
Global parameters can also vary in time. For instance, the external environment can exhibit wet–dry cycles [180] or temperature oscillations, which need to be compensated in circadian rhythms [181]. Alternatively, cells can generate oscillations internally to affect condensates [182]. These oscillatory drives can in particular influence chemical reactions controlled by droplets [183]. Besides structured variations of global parameters, cells are also subject to fluctuations, e.g., from gene expression. In the simple case of binary mixtures, these fluctuations can be fully buffered by phase separation since the equilibrium concentrations are independent of the total amount of material [184]. Buffering typically still works in multi-component mixtures, although the effect might be a bit weaker [185, 186].

### v2

@@ 2.1.4  Interfaces and phase morphologies
where $\bar{\mu}_{\mathrm{eq}}$ is the exchange chemical potential and $\Pi_{\mathrm{eq}}=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{out}}-f(\phi^{(0)}_{\mathrm{out}})=\frac{\bar{\mu}_{\mathrm{eq}}}{\nu}\phi^{(0)}_{\mathrm{in}}-f(\phi^{(0)}_{\mathrm{in}})$ denotes the associated osmotic pressure; see appendix C. The second line assumes a monotonic profile $\phi_{\mathrm{eq}}(x)$, which may not be valid in complex mixtures [128]. However, $\phi_{\mathrm{eq}}(x)$ is always monotonic in binary mixtures, and in the symmetric case $\Omega=1$, we find that
@@
The thin-interface approximation simplifies the analysis of equilibrium states in binary liquids: material conservation fixes the total volume of the dilute and dense phases, so their shape is the only degree of freedom. The total free energy thus comprises a constant term covering the bulk contributions and an interfacial term, $\gamma A$, where $A$ is the total interfacial area. Equilibrium states thus minimise $A$, which implies zero mean curvature of the interfaces [131]. Consequently, surface energies can be interpreted as a mechanical surface tension [132], explaining why the minority phase favours spherical shapes. Conversely, if both phases occupy similar volumes, bicontinuous structures, such as cylinders and flat lamella, can emerge; see figure 2.2. These structures predominate in a region of phase space that overlaps significantly with the spinodal region; see figure 2.1E,F. However, these two phenomena are distinct, so droplets can for instance emerge from spinodal decomposition. Bicontinuous structures are rarely observed in biological cells, and so we exclusively focus on droplets in this review.
@@ 2.2 Effective description of droplet dynamics
To focus on droplets, we next consider situations in which the dense phase forms spatially compact regions separated from the dilute phase by a thin interface. For simplicity, we focus on three-dimensional droplets (key results for two dimensions appear in appendix D) without reactions (which are added in section 5).
@@ 2.2.1  Local equilibrium at the interface
where $V_{\mathrm{sys}}$ is the system volume. For simplicity, we consider a spherical droplet of radius $R$ (implying volume $V=\frac{4\pi}{3}R^{3}$ and surface area $A=4\pi R^{2}$) with a surface tension $\gamma$ that is independent of the interface curvature, although this approximation can in principle be lifted [133]. Material conservation implies
@@
where $\bar{\phi}$ is the average component fraction in the system. Minimising $F$ using this constraint and fixed $V_{\mathrm{sys}}$ results in the coexistence conditions [23]
$\bar{\mu}(\phi_{\mathrm{in}})$ $=\bar{\mu}(\phi_{\mathrm{out}})\;,$ (2.30a)
@@
where $d$ is the dimension of space. These equations reduce to equations (2.5) if surface tension is negligible, $\gamma=0$. However, mechanical equilibrium as expressed in equation (2.30b) reveals that $\gamma$ generates an extra pressure $P_{\gamma}=(d-1)\gamma/R$, which is known as the Laplace pressure. Equations (2.30) are solved by the coexisting equilibrium fractions $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$, which depend on the droplet radius $R$ and only coincide with $\phi^{(0)}_{\mathrm{in}}$ and $\phi^{(0)}_{\mathrm{out}}$ for flat interfaces ($R\rightarrow\infty$). A pressure increase $\Delta P$ inside droplets generally implies larger chemical potentials, and thus larger $\phi^{\mathrm{eq}}_{\mathrm{in}}$ and $\phi^{\mathrm{eq}}_{\mathrm{out}}$. In the simple case of an incompressible droplet originating from strong phase separation, the fraction inside will hardly change, $\phi^{\mathrm{eq}}_{\mathrm{in}}\approx\phi^{(0)}_{\mathrm{in}}$, whereas the volume fraction outside approximately reads [134]
@@ 2.2.2  Classical nucleation theory
$J_{\mathrm{nucl}}$ $=J_{0}\exp\left(-\frac{\Delta\!F_{\mathrm{nucl}}}{k_{\mathrm{B}}T}\right)\;,$ (2.35)
@@ 2.2.3  Growth of a single droplet
Once a droplet has nucleated, it can grow by taking up material from its surroundings or by creating material inside itself via chemical reactions. For simplicity, we postpone the discussion of reactions to section 5 and first focus on the case without reactions ($s=0$). Growth of droplets of any shape can be described by the normal velocity $v_{\perp}$ of the droplet interface, which is given by [139, 23]
@@
where $\boldsymbol{n}$ is the normal vector to the interface. Here, $\boldsymbol{j}_{\mathrm{in}}$ and $\boldsymbol{j}_{\mathrm{out}}$ denote the fluxes directly inside and outside the interface, and so their difference is related to the net accumulation of material. The fluxes follow from the Cahn–Hilliard equation (2.23), but the thin-interface limit also provides a useful simplification since $\phi$ typically does not vary much either within or outside droplets. Consequently, we linearise equation (2.23) around the compositions $\phi^{(0)}_{\mathrm{in}}$ and $\phi^{(0)}_{\mathrm{out}}$ to obtain a simple diffusion equation
@@
where we have dropped the fourth-order derivative of $\phi$ and introduced the diffusivity
@@
As a simple example, we consider a spherical droplet embedded in a large system. Without chemical reactions, the droplet is essentially homogeneous, implying $\boldsymbol{j}_{\mathrm{in}}=0$. Conversely, the fraction $\phi_{\infty}$ far away from the droplet might deviate from its equilibrium fraction $\phi^{\mathrm{eq}}_{\mathrm{out}}$, resulting in diffusive fluxes around the droplet. Solving equation (2.37) at stationary state in three dimensions, we find ${\boldsymbol{j}_{\mathrm{out}}=(\phi^{\mathrm{eq}}_{\mathrm{out}}-\phi_{\infty})D_{\mathrm{out}}R^{-1}}$. The droplet growth rate following from equation (2.36) then reads [140]
@@
where we have neglected the small effects of surface tension in the denominator. Consequently, droplets grow spontaneously when the surrounding is supersaturated, $\phi_{\infty}>\phi^{\mathrm{eq}}_{\mathrm{out}}$. Moreover, droplets exhibit a drift when the surrounding concentration field is anisotropic [23],
@@
where the gradients are evaluated at the droplet position $\boldsymbol{x}$, and $d$ is the dimensionality of the system. Here, the anisotropy can originate from either concentration variations in the surrounding field $\phi_{\infty}$ or the equilibrium concentration itself, e.g., due to gradients in additional, regulating components [141, 110] or mechanical properties [142, 143].
Drift caused by anisotropic fields is often dominated by other processes that affect droplets’ positions. For instance, droplets in cells can be actively moved by molecular motors [144] and hydrodynamic flows [110]. Moreover, all droplets exhibit Brownian motion characterised by a diffusion constant [145]
@@
where $\eta$ denotes the viscosity of the surrounding liquid, which can depend on the droplet size in complex liquids such as the cytosol [146]. Equation (2.40) illustrates that droplets drift up supersaturation gradients, but Brownian motion caused by thermal fluctuations [145], hydrodynamic effects (discussed in section 3.3) and active transport may dominate their trajectory.
@@
which gives an explicit equation for $V$ in the limit $\phi^{\mathrm{eq}}_{\mathrm{out}}\approx\phi^{(0)}_{\mathrm{out}}$. This equation illustrates that droplet size is controlled by material conservation (equation 2.29), and that no droplets exist for volume fractions left of the binodal line ($\bar{\phi}<\phi^{(0)}_{\mathrm{out}}$).
@@ 2.2.4  Dynamics of many droplets
The dynamical equations derived above for individual droplets also allow us to discuss the interaction of many droplets as long as they are well-separated. We thus consider a collection of ${N_{\mathrm{d}}}$ droplets described by their positions $\boldsymbol{x}_{n}$ and radii $R_{n}$ for $n=1,\ldots,{N_{\mathrm{d}}}$. These droplets are embedded in a common dilute phase, whose composition is described by a field $\phi(\boldsymbol{x})$. In the typical situation where droplets occupy only a small fraction of the entire system, it is convenient to assume that $\phi(\boldsymbol{x})$ is defined across the entire system, and that droplets provide localised perturbations [140] to this field. In this case, we describe the dynamics of the dilute phase by
@@
where the first term accounts for diffusive transport, while the second term accounts for the material exchange with droplets of volume $V_{n}=\frac{4\pi}{3}R_{n}^{3}$ localised at $\boldsymbol{x}_{n}$. The growth and drift dynamics of the droplets follow from equations (2.39) and (2.40),
$\frac{\text{d}V_{n}}{\text{d}t}$ $=\frac{4\pi R_{n}D_{\mathrm{out}}}{\phi^{(0)}_{\mathrm{in}}-\phi^{(0)}_{\mathrm{out}}}\Bigl{(}\phi(\boldsymbol{x}_{n})-\phi^{\mathrm{eq}}_{\mathrm{out}}(R_{n})\Bigr{)}\;,$ (2.44a)
@@
Equations (2.43)–(2.44) describe how droplets exchange material via diffusion through the dilute phase. In particular, droplets grow by taking up material from the dilute phase when the fraction $\phi$ in their surrounding is larger than the saturation concentration $\phi^{\mathrm{eq}}_{\mathrm{out}}$. Equation (2.32) shows that $\phi^{\mathrm{eq}}_{\mathrm{out}}$ decreases with the droplet radius $R$, implying that larger droplets grow at the expense of smaller droplets. The resulting coarsening process, where the droplet count decreases while the mean size increases, is known as Ostwald ripening [147, 148]. In the limit where the diffusion of material between droplets is faster than their growth ($D_{\mathrm{out}}\gg L\,\text{d}R/\text{d}t$, where $L$ is the typical droplet separation), heterogeneities in the dilute phase equilibrate quickly, so that all droplets experience a similar fraction $\phi(\boldsymbol{x})\approx\phi_{\infty}$. In this case, equation (2.32) implies that droplets grow when their radius is larger than the critical radius
@@
and so $\langle V\rangle\propto t$ and $\langle{N_{\mathrm{d}}}\rangle\propto t^{-1}$ in $d=3$ dimensions.
Droplets also diffuse and coalesce, leading to fewer, larger droplets over time. Interestingly, this coarsening process exhibits the same scaling law as Ostwald ripening [151]. To see this, assume that coalescence roughly doubles the droplet volume, ${\Delta\!V\propto R^{3}}$, and the time between collisions scales as $\Delta t\propto L^{2}/D_{\mathrm{drop}}$ where $L\propto R$ is the average droplet separation and equation (2.41) implies $D_{\mathrm{drop}}\propto R^{-1}$. Taken together, $\Delta\!V/\Delta t\propto 1$, so the average volume grows linearly in time, similarly to Ostwald ripening. Although both coarsening processes obey the same scaling laws, the pre-factors can be very different and depend on the particular details of the system [23].
Droplet coarsening is generally driven by the thermodynamic tendency to reduce interfacial area. However, the coarsening dynamics depend on details of kinetic processes [152], so observed exponents vary with composition and the strength of phase separation [106, 153]. For instance, coarsening can be slowed down by trapped species [154], charges [155], sub-diffusion in the dilute phase [156], and elastic inhomogeneities [157]. In contrast, coarsening can be accelerated when it is conversion-limited [150, 156], when mechanical relaxation is relevant [158], and when hydrodynamic effects are significant [159, 160], e.g., via coalescence-induced coalescence [161, 162]. In biological cells, coarsening is further complicated by active production and degradation of material, which we consider in section 5. It is currently unclear what governs the size distribution of droplets in cells, although there is evidence that a competition between coalescence and nucleation is crucial [163] and that the material properties of the surrounding are important [164]. Taken together, this shows that cells can use many physical processes to control droplet size.
@@ 2.3.3  Temporal variations
Global parameters can also vary in time. For instance, the external environment can exhibit wet–dry cycles [181] or temperature oscillations, which need to be compensated in circadian rhythms [182]. Alternatively, cells can generate oscillations internally to affect condensates [183]. These oscillatory drives can in particular influence chemical reactions controlled by droplets [184]. Besides structured variations of global parameters, cells are also subject to fluctuations, e.g., from gene expression. In the simple case of binary mixtures, these fluctuations can be buffered by phase separation since the equilibrium concentrations are independent of the total amount of material [185]. Buffering effects persist in multi-component mixtures, although concentration buffering may be weaker [186, 187]. On the other hand, volume buffering is also then possible [188].